We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror.
The power was reduced via the PBS we introduced in Entry 7295.
While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent.
I stopped the regular MC autolocker and told the crontab to startup the low power Mc autolocker on op340m. Also, since we now have the new MC2 transmission setup, the power that gets to the 'regular' MC trans PD is lower, so I've lowered the lock threshold to 50 counts, from 100 counts.